\begin{thebibliography}{76}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Abdelwahap, Elfarash, and
  Eltanboly(2021)}]{abdelwahap2021applications}
Abdelwahap, M.; Elfarash, M.; and Eltanboly, A. 2021.
\newblock Applications Of Natural Language Processing In Healthcare Systems.
\newblock In \emph{The International Undergraduate Research Conference},
  volume~5, 111--115. The Military Technical College.

\bibitem[{Achilonu et~al.(2021)Achilonu, Fabian, Bebington, Singh, Nimako,
  Eijkemans, and Musenge}]{achilonu2021predicting}
Achilonu, O.~J.; Fabian, J.; Bebington, B.; Singh, E.; Nimako, G.; Eijkemans,
  M.; and Musenge, E. 2021.
\newblock Predicting colorectal cancer recurrence and patient survival using
  supervised machine learning approach: a South African population-based study.
\newblock \emph{Frontiers in Public Health}, 9.

\bibitem[{Adeoye et~al.(2022)Adeoye, Akinshipo, Thomson, and
  Su}]{adeoye2022artificial}
Adeoye, J.; Akinshipo, A.; Thomson, P.; and Su, Y.-X. 2022.
\newblock Artificial intelligence-based prediction for cancer-related outcomes
  in Africa: Status and potential refinements.
\newblock \emph{Journal of Global Health}, 12.

\bibitem[{Ahishakiye et~al.(2020)Ahishakiye, Wario, Mwangi, and
  Taremwa}]{ahishakiye2020prediction}
Ahishakiye, E.; Wario, R.; Mwangi, W.; and Taremwa, D. 2020.
\newblock Prediction of cervical cancer basing on risk factors using ensemble
  learning.
\newblock In \emph{2020 IST-Africa Conference (IST-Africa)}, 1--12. IEEE.

\bibitem[{Anantpadma et~al.(2019)Anantpadma, Lane, Zorn, Lingerfelt, Clark,
  Freundlich, Davey, Madrid, and Ekins}]{anantpadma2019ebola}
Anantpadma, M.; Lane, T.; Zorn, K.~M.; Lingerfelt, M.~A.; Clark, A.~M.;
  Freundlich, J.~S.; Davey, R.~A.; Madrid, P.~B.; and Ekins, S. 2019.
\newblock Ebola virus Bayesian machine learning models enable new in vitro
  leads.
\newblock \emph{ACS omega}, 4(1): 2353--2361.

\bibitem[{Ardila et~al.(2019)Ardila, Kiraly, Bharadwaj, Choi, Reicher, Peng,
  Tse, Etemadi, Ye, Corrado et~al.}]{ardila2019end}
Ardila, D.; Kiraly, A.~P.; Bharadwaj, S.; Choi, B.; Reicher, J.~J.; Peng, L.;
  Tse, D.; Etemadi, M.; Ye, W.; Corrado, G.; et~al. 2019.
\newblock End-to-end lung cancer screening with three-dimensional deep learning
  on low-dose chest computed tomography.
\newblock \emph{Nature medicine}, 25(6): 954--961.

\bibitem[{Balzer et~al.(2020)Balzer, Havlir, Kamya, Chamie, Charlebois, Clark,
  Koss, Kwarisiima, Ayieko, Sang et~al.}]{balzer2020machine}
Balzer, L.~B.; Havlir, D.~V.; Kamya, M.~R.; Chamie, G.; Charlebois, E.~D.;
  Clark, T.~D.; Koss, C.~A.; Kwarisiima, D.; Ayieko, J.; Sang, N.; et~al. 2020.
\newblock Machine learning to identify persons at high-risk of human
  immunodeficiency virus acquisition in rural Kenya and Uganda.
\newblock \emph{Clinical Infectious Diseases}, 71(9): 2326--2333.

\bibitem[{Banerjee et~al.(2021)Banerjee, Bhimireddy, Burns, Celi, Chen, Correa,
  Dullerud, Ghassemi, Huang, Kuo et~al.}]{banerjee2021reading}
Banerjee, I.; Bhimireddy, A.~R.; Burns, J.~L.; Celi, L.~A.; Chen, L.-C.;
  Correa, R.; Dullerud, N.; Ghassemi, M.; Huang, S.-C.; Kuo, P.-C.; et~al.
  2021.
\newblock Reading Race: AI Recognises Patient's Racial Identity In Medical
  Images.
\newblock \emph{arXiv preprint arXiv:2107.10356}.

\bibitem[{Bates(2019)}]{bates2019health}
Bates, M. 2019.
\newblock Health care chatbots are here to help.
\newblock \emph{IEEE pulse}, 10(3): 12--14.

\bibitem[{Bauer and Schedl(2019)}]{bauer2019global}
Bauer, C.; and Schedl, M. 2019.
\newblock Global and country-specific mainstreaminess measures: Definitions,
  analysis, and usage for improving personalized music recommendation systems.
\newblock \emph{PloS one}, 14(6): e0217389.

\bibitem[{Beigbeder(1991)}]{beigbeder1991role}
Beigbeder, Y. 1991.
\newblock \emph{The role and statuts of International Humanitarian Volunteers
  and Organizations: The right and duty to humanitarian assistance}, volume~12.
\newblock Martinus Nijhoff Publishers.

\bibitem[{Bharti et~al.(2020)Bharti, Bajaj, Batra, Lalit, Lalit, and
  Gangwani}]{bharti2020medbot}
Bharti, U.; Bajaj, D.; Batra, H.; Lalit, S.; Lalit, S.; and Gangwani, A. 2020.
\newblock Medbot: Conversational artificial intelligence powered chatbot for
  delivering tele-health after covid-19.
\newblock In \emph{2020 5th international conference on communication and
  electronics systems (ICCES)}, 870--875. IEEE.

\bibitem[{Bohr and Memarzadeh(2020)}]{bohr2020rise}
Bohr, A.; and Memarzadeh, K. 2020.
\newblock The rise of artificial intelligence in healthcare applications.
\newblock In \emph{Artificial Intelligence in healthcare}, 25--60. Elsevier.

\bibitem[{Chan et~al.(2019)Chan, Shan, Dahoun, Vogel, and
  Yuan}]{chan2019advancing}
Chan, H.~S.; Shan, H.; Dahoun, T.; Vogel, H.; and Yuan, S. 2019.
\newblock Advancing drug discovery via artificial intelligence.
\newblock \emph{Trends in pharmacological sciences}, 40(8): 592--604.

\bibitem[{Choi et~al.(2016)Choi, Bahadori, Schuetz, Stewart, and
  Sun}]{choi2016doctor}
Choi, E.; Bahadori, M.~T.; Schuetz, A.; Stewart, W.~F.; and Sun, J. 2016.
\newblock Doctor ai: Predicting clinical events via recurrent neural networks.
\newblock In \emph{Machine learning for healthcare conference}, 301--318. PMLR.

\bibitem[{Colubri et~al.(2016)Colubri, Silver, Fradet, Retzepi, Fry, and
  Sabeti}]{colubri2016transforming}
Colubri, A.; Silver, T.; Fradet, T.; Retzepi, K.; Fry, B.; and Sabeti, P. 2016.
\newblock Transforming clinical data into actionable prognosis models:
  machine-learning framework and field-deployable app to predict outcome of
  Ebola patients.
\newblock \emph{PLoS neglected tropical diseases}, 10(3): e0004549.

\bibitem[{David et~al.(2020)David, Thakkar, Mercado, and
  Engkvist}]{david2020molecular}
David, L.; Thakkar, A.; Mercado, R.; and Engkvist, O. 2020.
\newblock Molecular representations in AI-driven drug discovery: a review and
  practical guide.
\newblock \emph{Journal of Cheminformatics}, 12(1): 1--22.

\bibitem[{Fadhil(2018)}]{fadhil2018beyond}
Fadhil, A. 2018.
\newblock Beyond patient monitoring: Conversational agents role in telemedicine
  \& healthcare support for home-living elderly individuals.
\newblock \emph{arXiv preprint arXiv:1803.06000}.

\bibitem[{Fathi-Torbaghan and Meyer(1994)}]{fathi1994medusa}
Fathi-Torbaghan, M.; and Meyer, D. 1994.
\newblock MEDUSA: a fuzzy expert system for medical diagnosis of acute
  abdominal pain.
\newblock \emph{Methods of information in medicine}, 33(05): 522--529.

\bibitem[{Feigenbaum(1981)}]{feigenbaum1981expert}
Feigenbaum, E.~A. 1981.
\newblock Expert systems in the 1980s.
\newblock \emph{State of the art report on machine intelligence. Maidenhead:
  Pergamon-Infotech}.

\bibitem[{Fihavango, Mohsini, and Mselle(2021)}]{fihavango2021using}
Fihavango, T.; Mohsini, M.~H.; and Mselle, L.~J. 2021.
\newblock Using data mining techniques to predict obstetric fistula in
  Tanzania: A case of CCBRT.
\newblock In \emph{Handbook of Research on Nurturing Industrial Economy for
  Africa’s Development}, 169--191. IGI Global.

\bibitem[{Fleming, Van~der Merwe, and McFerren(2007)}]{fleming2007fuzzy}
Fleming, G.; Van~der Merwe, M.; and McFerren, G. 2007.
\newblock Fuzzy expert systems and GIS for cholera health risk prediction in
  southern Africa.
\newblock \emph{Environmental Modelling \& Software}, 22(4): 442--448.

\bibitem[{Fleming(2018)}]{fleming2018artificial}
Fleming, N. 2018.
\newblock How artificial intelligence is changing drug discovery.
\newblock \emph{Nature}, 557(7706): S55--S55.

\bibitem[{Garfinkel(1999)}]{garfinkel1999architects}
Garfinkel, S. 1999.
\newblock \emph{Architects of the information society: 35 years of the
  Laboratory for Computer Science at MIT}.
\newblock MIT press.

\bibitem[{Gibbs(2014)}]{gibbs2014evolution}
Gibbs, E. P.~J. 2014.
\newblock The evolution of One Health: a decade of progress and challenges for
  the future.
\newblock \emph{Veterinary Record}, 174(4): 85--91.

\bibitem[{Gonsalves et~al.(2019)Gonsalves, Thabtah, Mohammad, and
  Singh}]{gonsalves2019prediction}
Gonsalves, A.~H.; Thabtah, F.; Mohammad, R. M.~A.; and Singh, G. 2019.
\newblock Prediction of coronary heart disease using machine learning: an
  experimental analysis.
\newblock In \emph{Proceedings of the 2019 3rd International Conference on Deep
  Learning Technologies}, 51--56.

\bibitem[{Guo and Li(2018)}]{guo2018application}
Guo, J.; and Li, B. 2018.
\newblock The application of medical artificial intelligence technology in
  rural areas of developing countries.
\newblock \emph{Health equity}, 2(1): 174--181.

\bibitem[{Gupta and Katarya(2020)}]{gupta2020social}
Gupta, A.; and Katarya, R. 2020.
\newblock Social media based surveillance systems for healthcare using machine
  learning: a systematic review.
\newblock \emph{Journal of biomedical informatics}, 108: 103500.

\bibitem[{Gwagwa et~al.(2020)Gwagwa, Kraemer-Mbula, Rizk, Rutenberg, and
  De~Beer}]{gwagwa2020artificial}
Gwagwa, A.; Kraemer-Mbula, E.; Rizk, N.; Rutenberg, I.; and De~Beer, J. 2020.
\newblock Artificial Intelligence (AI) deployments in Africa: benefits,
  challenges and policy dimensions.
\newblock \emph{The African Journal of Information and Communication}, 26:
  1--28.

\bibitem[{Hagerty and Rubinov(2019)}]{hagerty2019global}
Hagerty, A.; and Rubinov, I. 2019.
\newblock Global AI ethics: a review of the social impacts and ethical
  implications of artificial intelligence.
\newblock \emph{arXiv preprint arXiv:1907.07892}.

\bibitem[{Harmon et~al.(2020)Harmon, Sanford, Xu, Turkbey, Roth, Xu, Yang,
  Myronenko, Anderson, Amalou et~al.}]{harmon2020artificial}
Harmon, S.~A.; Sanford, T.~H.; Xu, S.; Turkbey, E.~B.; Roth, H.; Xu, Z.; Yang,
  D.; Myronenko, A.; Anderson, V.; Amalou, A.; et~al. 2020.
\newblock Artificial intelligence for the detection of COVID-19 pneumonia on
  chest CT using multinational datasets.
\newblock \emph{Nature communications}, 11(1): 1--7.

\bibitem[{Holbrook et~al.(2021)Holbrook, Clark, Patel, Qi, Bassil, Mowery, and
  Badea}]{holbrook2021detection}
Holbrook, M.~D.; Clark, D.~P.; Patel, R.; Qi, Y.; Bassil, A.~M.; Mowery, Y.~M.;
  and Badea, C.~T. 2021.
\newblock Detection of Lung Nodules in Micro-CT Imaging Using Deep Learning.
\newblock \emph{Tomography}, 7(3): 358--372.

\bibitem[{Howe, Gardner, and Kadel(2006)}]{howe2006perioperative}
Howe, C.~R.; Gardner, G.~C.; and Kadel, N.~J. 2006.
\newblock Perioperative medication management for the patient with rheumatoid
  arthritis.
\newblock \emph{JAAOS-Journal of the American Academy of Orthopaedic Surgeons},
  14(9): 544--551.

\bibitem[{Hu et~al.(2018)Hu, Tang, Wang, Zhang, Zhang, and Sun}]{hu2018deep}
Hu, Z.; Tang, J.; Wang, Z.; Zhang, K.; Zhang, L.; and Sun, Q. 2018.
\newblock Deep learning for image-based cancer detection and diagnosis- A
  survey.
\newblock \emph{Pattern Recognition}, 83: 134--149.

\bibitem[{Ibrahim, Tulay, and Abdullahi(2022)}]{ibrahim2022multi}
Ibrahim, Z.; Tulay, P.; and Abdullahi, J. 2022.
\newblock Multi-region machine learning-based novel ensemble approaches for
  predicting COVID-19 pandemic in Africa.
\newblock \emph{Environmental Science and Pollution Research}, 1--23.

\bibitem[{Kamulegeya et~al.(2019)Kamulegeya, Okello, Bwanika, Musinguzi,
  Lubega, Rusoke, Nassiwa, and B{\"o}rve}]{kamulegeya2019using}
Kamulegeya, L.~H.; Okello, M.; Bwanika, J.~M.; Musinguzi, D.; Lubega, W.;
  Rusoke, D.; Nassiwa, F.; and B{\"o}rve, A. 2019.
\newblock Using artificial intelligence on dermatology conditions in Uganda: A
  case for diversity in training data sets for machine learning.
\newblock \emph{BioRxiv}, 826057.

\bibitem[{Kaul, Enslin, and Gross(2020)}]{kaul2020history}
Kaul, V.; Enslin, S.; and Gross, S.~A. 2020.
\newblock History of artificial intelligence in medicine.
\newblock \emph{Gastrointestinal endoscopy}, 92(4): 807--812.

\bibitem[{Lagomarsino et~al.(2012)Lagomarsino, Garabrant, Adyas, Muga, and
  Otoo}]{lagomarsino2012moving}
Lagomarsino, G.; Garabrant, A.; Adyas, A.; Muga, R.; and Otoo, N. 2012.
\newblock Moving towards universal health coverage: health insurance reforms in
  nine developing countries in Africa and Asia.
\newblock \emph{The Lancet}, 380(9845): 933--943.

\bibitem[{Lakshmanaprabu et~al.(2019)Lakshmanaprabu, Mohanty, Shankar,
  Arunkumar, and Ramirez}]{lakshmanaprabu2019optimal}
Lakshmanaprabu, S.; Mohanty, S.~N.; Shankar, K.; Arunkumar, N.; and Ramirez, G.
  2019.
\newblock Optimal deep learning model for classification of lung cancer on CT
  images.
\newblock \emph{Future Generation Computer Systems}, 92: 374--382.

\bibitem[{Lane et~al.(2019)Lane, Anantpadma, Freundlich, Davey, Madrid, and
  Ekins}]{lane2019natural}
Lane, T.; Anantpadma, M.; Freundlich, J.~S.; Davey, R.~A.; Madrid, P.~B.; and
  Ekins, S. 2019.
\newblock The natural product eugenol is an inhibitor of the ebola virus in
  vitro.
\newblock \emph{Pharmaceutical research}, 36(7): 1--6.

\bibitem[{Lavecchia(2019)}]{lavecchia2019deep}
Lavecchia, A. 2019.
\newblock Deep learning in drug discovery: opportunities, challenges and future
  prospects.
\newblock \emph{Drug discovery today}, 24(10): 2017--2032.

\bibitem[{Lee and Wang(2010)}]{lee2010fuzzy}
Lee, C.-S.; and Wang, M.-H. 2010.
\newblock A fuzzy expert system for diabetes decision support application.
\newblock \emph{IEEE Transactions on Systems, Man, and Cybernetics, Part B
  (Cybernetics)}, 41(1): 139--153.

\bibitem[{Lee et~al.(2018)Lee, Ragguett, Mansur, Boutilier, Rosenblat,
  Trevizol, Brietzke, Lin, Pan, Subramaniapillai et~al.}]{lee2018applications}
Lee, Y.; Ragguett, R.-M.; Mansur, R.~B.; Boutilier, J.~J.; Rosenblat, J.~D.;
  Trevizol, A.; Brietzke, E.; Lin, K.; Pan, Z.; Subramaniapillai, M.; et~al.
  2018.
\newblock Applications of machine learning algorithms to predict therapeutic
  outcomes in depression: a meta-analysis and systematic review.
\newblock \emph{Journal of affective disorders}, 241: 519--532.

\bibitem[{Li et~al.(2020)Li, Li, Li, Alvarez-Napagao, Garcia-Gasulla, and
  Suzumura}]{li2020we}
Li, I.; Li, Y.; Li, T.; Alvarez-Napagao, S.; Garcia-Gasulla, D.; and Suzumura,
  T. 2020.
\newblock What are we depressed about when we talk about covid-19: Mental
  health analysis on tweets using natural language processing.
\newblock In \emph{International Conference on Innovative Techniques and
  Applications of Artificial Intelligence}, 358--370. Springer.

\bibitem[{Lin et~al.(2019)Lin, Li, Liu, Chen, Yang, Chen, Lin, Lai, Long, Wu
  et~al.}]{lin2019diagnostic}
Lin, H.; Li, R.; Liu, Z.; Chen, J.; Yang, Y.; Chen, H.; Lin, Z.; Lai, W.; Long,
  E.; Wu, X.; et~al. 2019.
\newblock Diagnostic efficacy and therapeutic decision-making capacity of an
  artificial intelligence platform for childhood cataracts in eye clinics: a
  multicentre randomized controlled trial.
\newblock \emph{EClinicalMedicine}, 9: 52--59.

\bibitem[{Liu et~al.(2020)Liu, Jain, Eng, Way, Lee, Bui, Kanada,
  de~Oliveira~Marinho, Gallegos, Gabriele et~al.}]{liu2020deep}
Liu, Y.; Jain, A.; Eng, C.; Way, D.~H.; Lee, K.; Bui, P.; Kanada, K.;
  de~Oliveira~Marinho, G.; Gallegos, J.; Gabriele, S.; et~al. 2020.
\newblock A deep learning system for differential diagnosis of skin diseases.
\newblock \emph{Nature medicine}, 26(6): 900--908.

\bibitem[{Loftus et~al.(2020)Loftus, Tighe, Filiberto, Efron, Brakenridge,
  Mohr, Rashidi, Upchurch, and Bihorac}]{loftus2020artificial}
Loftus, T.~J.; Tighe, P.~J.; Filiberto, A.~C.; Efron, P.~A.; Brakenridge,
  S.~C.; Mohr, A.~M.; Rashidi, P.; Upchurch, G.~R.; and Bihorac, A. 2020.
\newblock Artificial intelligence and surgical decision-making.
\newblock \emph{JAMA surgery}, 155(2): 148--158.

\bibitem[{Lysaght et~al.(2019)Lysaght, Lim, Xafis, and Ngiam}]{lysaght2019ai}
Lysaght, T.; Lim, H.~Y.; Xafis, V.; and Ngiam, K.~Y. 2019.
\newblock AI-assisted decision-making in healthcare.
\newblock \emph{Asian Bioethics Review}, 11(3): 299--314.

\bibitem[{Macaulay et~al.(2021)Macaulay, Aribisala, Akande, Akinnuwesi, and
  Olabanjo}]{macaulay2021breast}
Macaulay, B.~O.; Aribisala, B.~S.; Akande, S.~A.; Akinnuwesi, B.~A.; and
  Olabanjo, O.~A. 2021.
\newblock Breast cancer risk prediction in African women using Random Forest
  Classifier.
\newblock \emph{Cancer Treatment and Research Communications}, 28: 100396.

\bibitem[{Makaju et~al.(2018)Makaju, Prasad, Alsadoon, Singh, and
  Elchouemi}]{makaju2018lung}
Makaju, S.; Prasad, P.; Alsadoon, A.; Singh, A.; and Elchouemi, A. 2018.
\newblock Lung cancer detection using CT scan images.
\newblock \emph{Procedia Computer Science}, 125: 107--114.

\bibitem[{Marais and Petersen(2015)}]{marais2015health}
Marais, D.~L.; and Petersen, I. 2015.
\newblock Health system governance to support integrated mental health care in
  South Africa: challenges and opportunities.
\newblock \emph{International journal of mental health systems}, 9(1): 1--21.

\bibitem[{Masinde(2020)}]{masinde2020africa}
Masinde, M. 2020.
\newblock Africa's Malaria Epidemic Predictor: Application of Machine Learning
  on Malaria Incidence and Climate Data.
\newblock In \emph{Proceedings of the 2020 the 4th International Conference on
  Compute and Data Analysis}, 29--37.

\bibitem[{Mbunge, Muchemwa, and Batani(2022)}]{mbunge2022we}
Mbunge, E.; Muchemwa, B.; and Batani, J. 2022.
\newblock Are we there yet? Unbundling the potential adoption and integration
  of telemedicine to improve virtual healthcare services in African health
  systems.
\newblock \emph{Sensors International}, 3: 100152.

\bibitem[{Mburu, Lee, and Mbogho(2012)}]{6428572}
Mburu, C.; Lee, H.; and Mbogho, A. 2012.
\newblock Ehealth advisory expert system for HIV/AIDS patients in South Africa.
\newblock In \emph{7th International Conference on Appropriate Healthcare
  Technologies for Developing Countries}, 1--4.

\bibitem[{Mutai et~al.(2021)Mutai, McSharry, Ngaruye, and
  Musabanganji}]{mutai2021use}
Mutai, C.~K.; McSharry, P.~E.; Ngaruye, I.; and Musabanganji, E. 2021.
\newblock Use of machine learning techniques to identify HIV predictors for
  screening in sub-Saharan Africa.
\newblock \emph{BMC medical research methodology}, 21(1): 1--11.

\bibitem[{Naicker et~al.(2009)Naicker, Plange-Rhule, Tutt, and
  Eastwood}]{naicker2009shortage}
Naicker, S.; Plange-Rhule, J.; Tutt, R.~C.; and Eastwood, J.~B. 2009.
\newblock Shortage of healthcare workers in developing countries--Africa.
\newblock \emph{Ethnicity \& disease}, 19(1): 60.

\bibitem[{Nkiruka, Prasad, and Clement(2021)}]{nkiruka2021prediction}
Nkiruka, O.; Prasad, R.; and Clement, O. 2021.
\newblock Prediction of malaria incidence using climate variability and machine
  learning.
\newblock \emph{Informatics in Medicine Unlocked}, 22: 100508.

\bibitem[{Nutbeam(2000)}]{nutbeam2000health}
Nutbeam, D. 2000.
\newblock Health literacy as a public health goal: a challenge for contemporary
  health education and communication strategies into the 21st century.
\newblock \emph{Health promotion international}, 15(3): 259--267.

\bibitem[{Obermeyer et~al.(2019)Obermeyer, Powers, Vogeli, and
  Mullainathan}]{obermeyer2019dissecting}
Obermeyer, Z.; Powers, B.; Vogeli, C.; and Mullainathan, S. 2019.
\newblock Dissecting racial bias in an algorithm used to manage the health of
  populations.
\newblock \emph{Science}, 366(6464): 447--453.

\bibitem[{Oleribe et~al.(2019)Oleribe, Momoh, Uzochukwu, Mbofana, Adebiyi,
  Barbera, Williams, and Taylor-Robinson}]{oleribe2019identifying}
Oleribe, O.~O.; Momoh, J.; Uzochukwu, B.~S.; Mbofana, F.; Adebiyi, A.; Barbera,
  T.; Williams, R.; and Taylor-Robinson, S.~D. 2019.
\newblock Identifying key challenges facing healthcare systems in Africa and
  potential solutions.
\newblock \emph{International journal of general medicine}, 12: 395.

\bibitem[{Owoyemi et~al.(2020)Owoyemi, Owoyemi, Osiyemi, and
  Boyd}]{owoyemi2020artificial}
Owoyemi, A.; Owoyemi, J.; Osiyemi, A.; and Boyd, A. 2020.
\newblock Artificial intelligence for healthcare in Africa.
\newblock \emph{Frontiers in Digital Health}, 2: 6.

\bibitem[{Oyebode and Orji(2018)}]{oyebode2018likita}
Oyebode, O.; and Orji, R. 2018.
\newblock Likita: a medical chatbot to improve healthcare delivery in Africa.
\newblock \emph{HCI Across Borders (HCIxB)}.

\bibitem[{Oyebode and Orji(2019)}]{oyebode2019detecting}
Oyebode, O.; and Orji, R. 2019.
\newblock Detecting factors responsible for diabetes prevalence in Nigeria
  using social media and machine learning.
\newblock In \emph{2019 15th International Conference on Network and Service
  Management (CNSM)}, 1--4. IEEE.

\bibitem[{Paul et~al.(2021)Paul, Sanap, Shenoy, Kalyane, Kalia, and
  Tekade}]{paul2021artificial}
Paul, D.; Sanap, G.; Shenoy, S.; Kalyane, D.; Kalia, K.; and Tekade, R.~K.
  2021.
\newblock Artificial intelligence in drug discovery and development.
\newblock \emph{Drug discovery today}, 26(1): 80.

\bibitem[{Senders et~al.(2018)Senders, Staples, Karhade, Zaki, Gormley,
  Broekman, Smith, and Arnaout}]{senders2018machine}
Senders, J.~T.; Staples, P.~C.; Karhade, A.~V.; Zaki, M.~M.; Gormley, W.~B.;
  Broekman, M.~L.; Smith, T.~R.; and Arnaout, O. 2018.
\newblock Machine learning and neurosurgical outcome prediction: a systematic
  review.
\newblock \emph{World neurosurgery}, 109: 476--486.

\bibitem[{Simi~Margarat et~al.(2022)Simi~Margarat, Hemalatha, Mishra, Shaheen,
  Maheswari, Tamijeselvan, Pavan~Kumar, Banupriya, Ferede
  et~al.}]{simi2022early}
Simi~Margarat, G.; Hemalatha, G.; Mishra, A.; Shaheen, H.; Maheswari, K.;
  Tamijeselvan, S.; Pavan~Kumar, U.; Banupriya, V.; Ferede, A.~W.; et~al. 2022.
\newblock Early diagnosis of tuberculosis using deep learning approach for iot
  based healthcare applications.
\newblock \emph{Computational Intelligence and Neuroscience}, 2022.

\bibitem[{Smalley(2017)}]{smalley2017ai}
Smalley, E. 2017.
\newblock AI-powered drug discovery captures pharma interest.
\newblock \emph{Nature Biotechnology}, 35(7): 604--606.

\bibitem[{Smith, Roitberg, and Isayev(2018)}]{smith2018transforming}
Smith, J.~S.; Roitberg, A.~E.; and Isayev, O. 2018.
\newblock Transforming computational drug discovery with machine learning and
  AI.

\bibitem[{Stafford et~al.(2020)Stafford, Kellermann, Mossotto, Beattie,
  MacArthur, and Ennis}]{stafford2020systematic}
Stafford, I.; Kellermann, M.; Mossotto, E.; Beattie, R.~M.; MacArthur, B.~D.;
  and Ennis, S. 2020.
\newblock A systematic review of the applications of artificial intelligence
  and machine learning in autoimmune diseases.
\newblock \emph{NPJ digital medicine}, 3(1): 1--11.

\bibitem[{Stephen et~al.(2019)Stephen, Sain, Maduh, and
  Jeong}]{stephen2019efficient}
Stephen, O.; Sain, M.; Maduh, U.~J.; and Jeong, D.-U. 2019.
\newblock An efficient deep learning approach to pneumonia classification in
  healthcare.
\newblock \emph{Journal of healthcare engineering}, 2019.

\bibitem[{Taconet et~al.(2021)Taconet, Porciani, Soma, Mouline, Simard, Koffi,
  Pennetier, Dabir{\'e}, Mangeas, and Moiroux}]{taconet2021data}
Taconet, P.; Porciani, A.; Soma, D.~D.; Mouline, K.; Simard, F.; Koffi, A.~A.;
  Pennetier, C.; Dabir{\'e}, R.~K.; Mangeas, M.; and Moiroux, N. 2021.
\newblock Data-driven and interpretable machine-learning modeling to explore
  the fine-scale environmental determinants of malaria vectors biting rates in
  rural Burkina Faso.
\newblock \emph{Parasites \& Vectors}, 14(1): 1--23.

\bibitem[{Thompson et~al.(2017)Thompson, Sowunmi, Misra, Fernandez-Sanz,
  Crawford, and Soto}]{thompson2017expert}
Thompson, T.; Sowunmi, O.; Misra, S.; Fernandez-Sanz, L.; Crawford, B.; and
  Soto, R. 2017.
\newblock An expert system for the diagnosis of sexually transmitted
  diseases--ESSTD.
\newblock \emph{Journal of Intelligent \& Fuzzy Systems}, 33(4): 2007--2017.

\bibitem[{Tran et~al.(2019)Tran, Vu, Ha, Vuong, Ho, Vuong, La, Ho, Nghiem,
  Nguyen et~al.}]{tran2019global}
Tran, B.~X.; Vu, G.~T.; Ha, G.~H.; Vuong, Q.-H.; Ho, M.-T.; Vuong, T.-T.; La,
  V.-P.; Ho, M.-T.; Nghiem, K.-C.~P.; Nguyen, H. L.~T.; et~al. 2019.
\newblock Global evolution of research in artificial intelligence in health and
  medicine: a bibliometric study.
\newblock \emph{Journal of clinical medicine}, 8(3): 360.

\bibitem[{Vaishya et~al.(2020)Vaishya, Javaid, Khan, and
  Haleem}]{vaishya2020artificial}
Vaishya, R.; Javaid, M.; Khan, I.~H.; and Haleem, A. 2020.
\newblock Artificial Intelligence (AI) applications for COVID-19 pandemic.
\newblock \emph{Diabetes \& Metabolic Syndrome: Clinical Research \& Reviews},
  14(4): 337--339.

\bibitem[{Wahl et~al.(2018)Wahl, Cossy-Gantner, Germann, and
  Schwalbe}]{wahl2018artificial}
Wahl, B.; Cossy-Gantner, A.; Germann, S.; and Schwalbe, N.~R. 2018.
\newblock Artificial intelligence (AI) and global health: how can AI contribute
  to health in resource-poor settings?
\newblock \emph{BMJ global health}, 3(4): e000798.

\bibitem[{Wakunuma, Jiya, and Aliyu(2020)}]{wakunuma2020socio}
Wakunuma, K.; Jiya, T.; and Aliyu, S. 2020.
\newblock Socio-ethical implications of using AI in accelerating SDG3 in Least
  Developed Countries.
\newblock \emph{Journal of Responsible Technology}, 4: 100006.

\end{thebibliography}

\def\year{2023}\relax
\documentclass[letterpaper]{article} % DO NOT CHANGE THIS
\usepackage{aaai23}  % DO NOT CHANGE THIS
\usepackage{times}  % DO NOT CHANGE THIS
\usepackage{helvet}  % DO NOT CHANGE THIS
\usepackage{courier}  % DO NOT CHANGE THIS
\usepackage[hyphens]{url}  % DO NOT CHANGE THIS
\usepackage{graphicx} % DO NOT CHANGE THIS
\usepackage{natbib}  % DO NOT CHANGE THIS AND DO NOT ADD ANY OPTIONS TO IT
\usepackage{caption} % DO NOT CHANGE THIS AND DO NOT ADD ANY OPTIONS TO IT
\DeclareCaptionStyle{ruled}{labelfont=normalfont,labelsep=colon,strut=off} % DO NOT CHANGE THIS
\usepackage{algorithm}
\usepackage{algorithmic}
\usepackage{blindtext}
\usepackage[utf8]{inputenc}
\DeclareUnicodeCharacter{2009}{\,}

\usepackage{newfloat}
\usepackage{listings}
\floatstyle{ruled}
\newfloat{listing}{tb}{lst}{}
\floatname{listing}{Listing}

\title{What We Know So Far: Artificial Intelligence in African Healthcare}
\author{
    Naome Etori,   %\thanks{With help from the AAAI Publications Committee.}\\
    Ebasa Temesgen, and
    Maria Gini \\
}
\affiliations{
    Department of Computer Science and Engineering\\
    University of Minnesota\\
    200 Union Street SE Minneapolis, MN 55455 \\
    \{etori001,temes021, gini\}@umn.edu
}

\usepackage{bibentry}
\begin{document}

\maketitle

\begin{abstract}

Healthcare in Africa is a complex issue influenced by many factors including poverty, lack of infrastructure, and inadequate funding. However, Artificial intelligence (AI) applied to healthcare, has the potential to transform healthcare in Africa by improving the accuracy and efficiency of diagnosis, enabling earlier detection of diseases, and supporting the delivery of personalized medicine. This paper reviews the current state of how AI Algorithms can be used to improve diagnostics, treatment, and diseases monitoring, as well as how AI can be used to improve access to healthcare in Africa as a low-resource setting and discusses some of the critical challenges and opportunities for its adoption. As such, there is a need for a well-coordinated effort by the governments, private sector, healthcare providers, and international organizations to create sustainable AI solutions that meet the unique needs of the African healthcare system.
\end{abstract}

\section{Introduction}
The application of AI in healthcare dates back to the 1950s. Researchers at the Massachusetts Institute of Technology (MIT) developed ``Project MAC'' a program to help analyze and interpret medical data, one of the first examples of AI in healthcare. MAC stood for ``machine-aided cognition'' \cite{garfinkel1999architects}.

In the 1960s, AI was used to create knowledge-based expert systems, that mimic the decision-making processes of human experts. The ``MYCIN'' system, developed at Stanford University to aid in diagnosing and treating infectious diseases, was an early example of an expert system in healthcare \cite{feigenbaum1981expert}.

During the 1980s and 1990s, AI-based Machine learning algorithms could learn and adapt to new data without being explicitly programmed. This led to the development of programs that could analyze medical images, such as X-rays, to help diagnose diseases and AI algorithms used for diagnosing, identifying, and analyzing public health threats \cite{kaul2020history}.

More recently, the improvements in  natural language processing (NLP) has been used to understand and interpret human language. NLP techniques have been applied in healthcare to extract and analyze data such as electronic health records (EHRs),medical images and assist with tasks (such as appointment scheduling and medication management). As a result of increases in computing power, the availability of large amounts of data, and the rapid development of big data analytics have enabled AI applications. Deep learning (DL) machine learning technology, which involves training artificial neural networks (ANN) on large datasets, has been a major driver of recent AI advances. This has significantly impacted modernizing the global healthcare system to improve diagnosis and clinical care accuracy and efficiency. A convolutional neural network (CNN) is a type of DL algorithm that simulates the behavior of interconnected neurons in the human brain \cite{kaul2020history}.

The recent surge in AI healthcare research in Africa indicates the potential of AI to improve patient outcomes and reduce the burden on the healthcare system. This paper discusses AI algorithms (such as expert systems, machine learning, deep learning, natural language processing, and image processing) and how they are applied in the African healthcare systems, including challenges in the application of AI systems in Africa as well as opportunities for AI adoption. 

\section{Problem Definition}
There has been a lot of research on AI in healthcare, but there has been minimal discussion about AI in African healthcare. AI researchers have also found that there is a lack of diverse representation of AI models in healthcare and particularly from low-resource languages. To fill this gap, we address two main research questions:

\textbf{RQ1:} \textit{ What role do AI algorithms play in African healthcare systems?} 

\textbf{RQ2:} \textit{How does a lack of resources impact AI implementation in Africa?} 

Our goal was to understand and explain the current integration and implementation of AI in African healthcare systems to inspire future AI research in African Healthcare.

\section{Related Work}
Many researchers have shown that AI has the potential to improve patient care and lower healthcare costs. Our review of the literature yielded the following topics for discussion.

\subsection{Recent AI Advances in Healthcare in Developed Countries }
AI has made significant advancements in the healthcare industry in developed countries in recent years. This is because AI algorithms can handle and learn from vast amount of healthcare data generated due to advancement of AI technologies. Deep learning, in particular has impacted how we view AI tools today and is the source of the recent advancement in AI applications \cite{bohr2020rise}.

Sophisticated AI algorithms have been applied in diagnosis and treatment planning to analyze medical images, such as X-rays and CT scans, to help doctors diagnose diseases \cite{liu2020deep,vaishya2020artificial,lin2019diagnostic,hu2018deep} and decide plan treatment. For example, \cite{lakshmanaprabu2019optimal,holbrook2021detection,makaju2018lung,ardila2019end} shows how an AI algorithm was able to detect lung cancer from CT scans with a very high degree of accuracy, comparable to that of an experienced radiologist. 

Big data has created opportunities for AI-based Machine learning (ML) algorithms, enabling them to learn from previous data to make accurate predictions about new data. This can assist healthcare providers in prioritizing their resources and intervening before a patient condition deteriorates\cite{senders2018machine,stafford2020systematic,lee2018applications,choi2016doctor}. 

Pharmaceutical companies use AI algorithms to identify new candidates for drug development and analyze the effects of different drugs \cite{david2020molecular,paul2021artificial,smith2018transforming,smalley2017ai,fleming2018artificial,chan2019advancing,lavecchia2019deep}.

AI-powered Virtual assistants help triage patients and provide them with information about their condition and treatment options \cite{bharti2020medbot,bates2019health,fadhil2018beyond}

The use of clinical decision support \cite{loftus2020artificial,lysaght2019ai} and medication management \cite{howe2006perioperative} have improved patient care while reducing the workload for healthcare professionals in hospitals.

\subsection{Recent AI Advances in African Healthcare}
AI has enormous potential to help in the prediction and prevention of outbreaks of infectious diseases. Data from the World Health Organization (WHO) \cite{nkiruka2021prediction} shows the relationship between malaria incidence and various climate variables, found that specific patterns in the data were indicative of increased malaria risk, and used machine learning algorithms to create a model for predicting malaria incidence based on these patterns. Research conducted by \cite{colubri2016transforming} used past Ebola outbreak data to train the ML algorithm to accurately predict the outcomes of Ebola patients with a high degree of accuracy which could be used in future outbreaks to help guide clinical decision-making and improve patient outcomes.

AI can improve the accuracy of diagnoses and treatment recommendations in Africa. Due to a lack of resources for optimal healthcare provision, trained physicians, and underfunded public healthcare facilities, ML models can be used to analyze medical images and provide diagnoses and chronic diseases such as cancer. \cite{adeoye2022artificial} shows the application and implementation phases of oncological AI tools in Africa using patient cohorts in Africa.  

\subsection{Status of AI in African Healthcare system}
The application of AI in the African healthcare system is still in its early stages. However, there is a growing interest and investment in the implementation of AI to improve different aspects of healthcare delivery in Africa. However, the implementation of AI in Africa confronts numerous challenges\cite{owoyemi2020artificial}. Many African countries, for example, still have limited access to reliable electricity and to high-speed internet, making it challenging to implement and use AI systems. The successful implementation of AI in health care depends on the availability and dependability of the infrastructure\cite{mbunge2022we}. 

For many years, African healthcare systems have suffered from man-made issues such as institutional, human resource, financial, technical, and political developments \cite{oleribe2019identifying}. Hence, most African countries cannot meet the fundamental requirements for effective healthcare systems. Ineffective service integration is linked to poor governance, and human resource challenges \cite{nutbeam2000health,marais2015health}. The lack of basic access to healthcare makes it even harder to implement AI solutions  \cite{tran2019global}.

Additionally, the scarcity of trained professionals with expertise in AI in Africa is a major concern; hence, many healthcare systems are struggling to meet the rising demand for services while also facing significant shortages of trained health workers and essential medicines \cite{naicker2009shortage,guo2018application}. As a result, most African countries rely heavily on developed nations for AI technologies to solve critical healthcare challenges due to inadequate funding \cite{guo2018application}. 

Ethical considerations around the use of AI in healthcare in Africa, including issues related to data privacy and the potential for biased decision-making could contribute to slow AI implementation in the African region \cite{hagerty2019global}.

\subsection{AI Gaps in Africa: Key Challenges}
African healthcare systems are lagging in implementing AI  mainly due to lack of resources. Implementing AI in healthcare can be expensive, and securing funding for such initiatives can be challenging in most African countries. Ensuring that AI models are sustainable in the long term is also an important consideration \cite{wakunuma2020socio}. 

\begin{itemize}
\item There is often a lack of high-quality data available for training machine learning models due to poor record-keeping and inadequate infrastructure for data collection and storage. For example, doctors and nurses still use hand written notes when seeing patients. This can make it difficult to develop accurate and effective AI systems for the African region. Prior studies have demonstrated disparate AI performance by race AI models, especially on African datasets. The training datasets may not be representative of the African population. If the datasets used to train the models are not diverse or representative enough, the models may not be able to generalize well to new, unseen data from Africa population \cite{banerjee2021reading,kamulegeya2019using,harmon2020artificial,obermeyer2019dissecting}.

\item Furthermore, the AI models may not have been specifically designed to handle the unique challenges present in African datasets. For example, African datasets may contain a greater variety of languages, dialects, and accents, and the models may not have been trained on data that reflects this linguistic diversity. In addition, there may be other cultural and societal factors that are unique to Africa that the models have not been designed to handle. According to estimates, 17\% of the world's languages, many of which are spoken in Africa, are "low-resource languages" in the digital realm \cite{gwagwa2020artificial}.

\item Despite these challenges, there have been some efforts to improve healthcare in Africa. For example, some African governments have increased funding for healthcare, and there have been efforts to train more healthcare professionals and build new healthcare facilities, as many developing countries are attempting to move more towards universal healthcare coverage \cite{lagomarsino2012moving}. In addition, some international organizations and charities have provided assistance to boost the delivery of healthcare in Africa \cite{gibbs2014evolution,beigbeder1991role}.

\end{itemize}

\section{Methodology}
This paper reviews the current state of AI in healthcare in Africa, including the challenges and opportunities for its adoption. A comprehensive search of the literature was conducted using several databases. The survey reviewed 30 journal papers obtained electronically through four scientific databases (Google Scholar, Scopus, IEEE, Pum Med, and Science Direct) searched using three sets of keywords: (1) Artificial Intelligence in Africa (2) Artificial Intelligence in Healthcare (3) African Healthcare Systems. We limited our search to articles that made algorithmic contributions and addressed diseases in their applications. We disqualified papers that only provided opinions, theories, surveys, or datasets. This yielded a total of 12 papers.

\section{Findings}
\subsection{RQ1: Role of AI Algorithms on African Healthcare }
Based on our findings, different algorithms have been used to address different healthcare,  such as Expert systems, machine learning, and deep learning. Although  expert systems have been extensively researched globally, little research has been conducted on its application in African healthcare systems. Expert systems can be beneficial, especially where scarcity of trained physicians and medical personnel exists, as is usually the case in African countries. Expert systems can help African healthcare systems diagnose patients and select treatment plans without extensively trained medical personnel, where a decision must be made quickly to save lives \cite{wahl2018artificial}. An example is where an expert system has been incorporated with fuzzy logic systems, to improve the diagnosis of chronic conditions like STDs, HIV/AIDS, cholera, abdominal pain, and diabetes decision support application in South Africa \cite{6428572, thompson2017expert, fathi1994medusa, fleming2007fuzzy, lee2010fuzzy}. 

Natural language processing (NLP) has been used in developing a Medical Chatbot to diagnose patients in their early stages of the disease or to use social media data for surveillance and monitoring of infectious disease outbreaks \cite{gupta2020social, abdelwahap2021applications}. 
The use of NLP in African healthcare is still at its infancy stage. \cite{li2020we} conducted mental health condition research following the outbreak of coronavirus(COVID-19), using Twitter data from Nigeria and South Africa. \cite{oyebode2018likita} demonstrated that Likita, a chatbot, could be used to diagnose common ailments and improve healthcare delivery in Africa.

Deep learning (DL) can process large amounts of data, such as images, and could potentially aid medical workers in decision-making, using an X-ray image to analyze multiple diseases. For instance, \cite{stephen2019efficient} used X-ray images to classify Pneumonia with a validation accuracy of 93.73\%. \cite{simi2022early} also used X-ray data to diagnose early stage Tuberculosis using a DL algorithm and achieved 99\% accuracy. \cite{holbrook2021detection} used three pre-trained DL models (faster R-CNN, single-shot multi-box detector (SSD), and RetinaNet) for the microscopic diagnosis of malaria parasites, malaria is one of the deadliest diseases in Sub-Sahara Africa in thick blood smears. The result found that a faster R-CNN has a higher accuracy than the other two models used in the study with an average precision of over 0.94. Hence this approach can improve the accuracy and efficiency of malaria diagnosis compared to traditional methods.

ML models are used to predict and classify chronic diseases in Africa. They are gaining popularity for their simplicity and the availability of data in a few African countries. Table 1 has summarized some of the research on applying AI to address different diseases in  African countries.

\subsection{RQ2: How does the lack of resources impact AI implementation in Africa?}
African healthcare systems suffer from the lack of data access due to inadequate resources, which is often a key ingredient in the development and training of AI systems. Lack of infrastructure and investment in data collection and storage has made it difficult for organizations to obtain the data they need to develop and train AI systems \cite{banerjee2021reading,kamulegeya2019using,harmon2020artificial,obermeyer2019dissecting}

%\subsection{Deep Learning}
%Deep learning (DL) is used to process large amounts of data, such as images, and could potentially aid medical workers in decision-making. Using an X-ray image to analyze multiple diseases. For instance, \cite{stephen2019efficient} used X-ray images to classify Pneumonia with a validation accuracy of 93.73\%. \cite{simi2022early} also used X-ray data to diagnose Tuberculosis at an early stage using a Deep learning algorithm and achieved 99\% accuracy. 

%\subsection{Machine Learning}
%Its simplicity of application and the availability of data in a few African countries has increased the research on Machine learning. Table 1 has summarized some of the research that has been conducted in the application of AI to address different types of diseases in different African countries.   

\begin{table*}[!htbp]
\caption{Application of AI and Machine Learning in Africa healthcare System}
\label{tab:table}
\begin{tabular}
{| p{0.1\linewidth}|p{0.15\linewidth}|p{0.15\linewidth} | p{0.15\linewidth} | p{0.15\linewidth} | p{0.16\linewidth} |}

\hline
\textbf{Disease} & \textbf{Country} & \textbf{Algorithm} & \textbf{Dataset} & \textbf{Accuracy} & \textbf{Reference} 

\\
\hline
HIV/AIDS & S.Saharan.A & XGBoost  &  PHIA & 90\% males, 92\% female & \cite{mutai2021use} \\
\hline
HIV/AIDS & Kenya, Uganda  &  - &Data from 16 communities  & 78\% & \cite{balzer2020machine} \\
\hline
Malaria & Burkina Faso  &  Random forest &georeferenced raster(SPOT)&84\% & \cite{taconet2021data} \\
\hline
Ebola & W.Africa & Central Bayesian & 842 molecules& 83\% & \cite{anantpadma2019ebola}
\\
\hline

Colorectal Cancer & S.Africa & Artificial NN & WDGMC CRC & 87.0\% & \cite{achilonu2021predicting}\\
\hline
Breast cancer & Nigeria & Random Forest &Surgical data(LASUTH) &  96.67\%  & \cite{macaulay2021breast} \\
\hline
Cervical cancer & African countries & Ensemble learning &  SMOTE & 87.21\% & \cite{ahishakiye2020prediction}\\ 

\hline
COVID-19 & African Countries& Ensemble  &  (WHO) database & MAD = 0.0073, MSE = 0.0002, R2 = 0.9616 & (Ibrahim et al. 2023)\\
\hline
Diabetes & Nigeria &  Binarized Naïve Bayes
(BNB) & Collected(nairaland) &  87.08\% & \cite{oyebode2019detecting}\\
\hline
Coronary Heart D & South Africa &  Naïve Bayes,SVM and Decision Tree & S.Africa-KEEL & higher than 70\% & 
 \cite{gonsalves2019prediction}\\ 
 
\hline
Obstetric Fistula & Tanzania & Logistic Regression& Collected from
CCBRT,Dar-es-Salaam & 86\% & (Fihavango et al. 2021) \\

\hline
Malaria & Sub-Saharan & Decision Tree & (WHO) datasets & 75.10\% & \cite{masinde2020africa} \\
\hline

\end{tabular}
\end{table*}

AI in African healthcare has generally been used in diseases mapping such as HIV. ML techniques identify HIV predictors for screening in sub-Saharan Africa \cite{mutai2021use}, malaria predictions \cite{nkiruka2021prediction}. Even though African healthcare systems strive to implement AI to assist with healthcare operations, it is still difficult to respond to public health emergencies such as disease outbreaks, resulting in increased mortality and morbidity.

\section{Discussion}
AI has been applied in various healthcare settings in Africa,  as shown in Table \ref{tab:table}. Hence has the potential to significantly improve Human Immunodeficiency Virus (HIV) care in Africa by providing more accurate diagnoses, predicting patient outcomes, and optimizing treatment programs. For example, \cite{bauer2019global} used DL to improve the accuracy of HIV diagnoses in sub-Saharan Africa, \cite{mutai2021use} applied ML approaches for building models in identifying HIV predictors as well as predicting persons at high risk of infection, the XGBoost algorithm significantly improved the identification of HIV positivity by f1 scoring mean of 90 and 92\% for males and females. respectively. Also, \cite{balzer2020machine} used  ML to Identify Persons at High-Risk of HIV Acquisition in rural Kenya and Uganda, and the result shows that ML improved efficiency by 78\%. 

Malaria has been (and still is) a major public health crisis in Africa, with an estimated 435 million cases and 1.3 million deaths yearly (WHO, 2021). Experts say malaria slows economic growth in Africa by up to 1.3 percent per year. AI can be used in the fight against malaria in Africa; for example, \cite{taconet2021data} used data-driven and interpretable ML modeling Algorithm (Random Forest) to explore malaria vector biting rates in rural Burkina Faso. The results identified several aspects of the bio-ecology of the main malaria vectors \cite{masinde2020africa} Decision trees algorithm and climate data produced accuracy results of 99\%.

Furthermore, AI has the potential to significantly improve the response to Ebola outbreaks in Africa by providing more accurate and timely predictions, faster diagnoses, and more efficient use of resources; for example, Bayesian (ML) Models were used to enable new in Vitro Leads\cite{anantpadma2019ebola,lane2019natural}.

Research has highlighted the potential benefits of using AI to improve cancer diagnosis and treatment in Africa. For example, in predicting colorectal cancer recurrence and patient survival using Artificial neural network(ANN), a supervised ML approach where scored the highest AUC-ROC for recurrence (87\%) and survival (82\%) in South Africa \cite{achilonu2021predicting}. Random forest (ML) models have been used in Breast cancer risk prediction among African women in Nigeria by \cite{macaulay2021breast}; the Chi-Square selected features gave the best performance with 98.33\% accuracy, 100\% sensitivity, 96.55\% specificity, and 98\% AUC. Also,\cite{ahishakiye2020prediction} predicted Cervical Cancer Based on risk factors using the Ensemble Learning model with an accuracy of 87.21\%.

Since the COVID-19 pandemic, the demand for use of AI in the African healthcare sector to aid in the pandemic response has increased. Many African countries have been conducting studies on this area; for example, \cite{ibrahim2022multi} conducted a study to predict daily COVID-19 cases spreading across the north, south, east, west, and central Africa regions and countries using  ML, including Artificial neural network(ANN), adaptive neuro-fuzzy inference system (ANFIS), support vector machine (SVM) and convent and the result shows high accuracy performance. 

According to the International Diabetes Federation (IDF), 24 million adults (ages 20-79) have diabetes, which accounts for 416,000 deaths in the IDF Africa Region in 2021. AI can potentially improve diabetes care in Africa by improving access to care and providing more personalized treatment options. For example, using social media and ML  \cite{oyebode2019detecting} detected factors responsible for Diabetes prevalence in Nigeria.

Management and treatment of coronary heart disease (CHD) in African healthcare systems have been a huge challenge. Recently, there has been a  growing recognition of the potential for AI to improve CHD disease treatment and care. For example, by using historical medical and ML models such as Naïve Bayes, SVM, and Decision Tree (DT) data to predict CHD  \cite{gonsalves2019prediction}.

WHO estimated more than 2 million young women live with untreated obstetric fistula (OF) in Asia and sub-Saharan Africa. Limited research on the use of AI in obstetric fistula disease management and treatment in Africa has been seen. For example, \cite{fihavango2021using} used Data Mining techniques to Predict Obstetric Fistula in Tanzania.

This survey did not cover the entire African continent. The results are promising, and the future of AI tools and algorithms to improve healthcare in the African continent look promising.

\section{Conclusions}
AI has the potential to predict and control diseases, expand and augment service delivery, and address several lingering social inequities by improving the accuracy and efficiency of diagnosis, enabling earlier detection of diseases, and supporting the delivery of personalized medicine. However, several challenges must be addressed to realize its full potential, including the lack of infrastructure, limited access to data, and the need for regulatory frameworks. African countries must build the necessary infrastructure to support technological advancements. This can include investing in high-speed internet, data centers, and cyber security systems. Governments can help promote the development and adoption of AI by establishing clear AI regulatory standards. Private sector, healthcare providers, partnership with international stakeholders can design a sustainable AI solutions that meet the unique needs of the African healthcare system

\section{Future Work}
AI and machine learning are critical to addressing health care inadequacies in African countries. AI models may perform poorly on African datasets, However, to improve the performance of AI models on African datasets, it will be necessary to create a more diverse and representative datasets as well as design models that are specifically tailored to deal with the unique challenges presented in African data.

% Use instead or the References section will not appear in your paper
\bibliography{aaai23}
% \section{Acknowledgments}
% AAAI is especially grateful to Peter Patel Schneider for his work in implementing the original aaai.sty file, liberally using the ideas of other style hackers, including Barbara Beeton. We also acknowledge with thanks the work of George Ferguson for his guide to using the style and BibTeX files.

\end{document}